OpenStudy (anonymous):

Find the slope of the graph of the equation: y = 5 - 3x

OpenStudy (luigi0210):

Put it into \(\Large y=mx+b\) and \(m\) will be your slope.

OpenStudy (anonymous):

so m would be 5

OpenStudy (anonymous):

or 3

OpenStudy (luigi0210):

\(\Large y=5-3x==> y=-3x+5\)

OpenStudy (luigi0210):

It's the whole coefficient, so "-3"

OpenStudy (anonymous):

oh ok so i have to solve the equation then ill find the answer

OpenStudy (luigi0210):

If you have the equation, yes :)

OpenStudy (anonymous):

ok thanks

OpenStudy (just_one_last_goodbye):

hmmm Great Job Luigi ^_^ no direct answer you deserve a medal OS should be like this helping the user step by step in the problems

OpenStudy (anonymous):

The coefficient of x will be your Slope if your given equation is in y = mx + c form..

OpenStudy (anonymous):

so would it be -1/3